# Link Prediction with Untrained Message Passing Layers

### v1

@@ Abstract
Message passing neural networks (MPNNs) operate on graphs by exchanging information between neigbouring nodes. MPNNs have been successfully applied to various node-, edge-, and graph-level tasks in areas like molecular science, computer vision, natural language processing, and combinatorial optimization. However, most MPNNs require training on large amounts of labeled data, which can be costly and time-consuming. In this work, we explore the use of various untrained message passing layers in graph neural networks, i.e. variants of popular message passing architecture where we remove all trainable parameters that are used to transform node features in the message passing step. Focusing on link prediction, we find that untrained message passing layers can lead to competitive and even superior performance compared to fully trained MPNNs, especially in the presence of high-dimensional features. We provide a theoretical analysis of untrained message passing by relating the inner products of features implicitly produced by untrained message passing layers to path-based topological node similarity measures. As such, untrained message passing architectures can be viewed as a highly efficient and interpretable approach to link prediction.
@@ 1 Introduction
Graph neural networks (GNNs) are a powerful class of machine learning models that can learn from graph-structured data, such as social networks, molecular graphs, and knowledge graphs. GNNs have emerged as an important tool in the machine learning landscape, due to their ability to model complex relationships and dependencies within data with applications in a variety of fields where data exhibits a complex topology that can be captured in a graph. This is shown by a multitude of studies, including Basu et al. (2022); Wang et al. (2022); Fu et al. (2022); Jaini et al. (2021); Puny et al. (2021), which highlight the versatility and adaptability of GNNs for machine learning tasks across a range of fields.
@@
One of the key concepts underlying GNNs is message passing (MP) Gilmer et al. (2017), which operates by propagating and aggregating information between nodes in the graph, using message and update functions possibly with learnable parameters. However, designing and training effective GNNs can be challenging, as they may suffer from issues such as over-smoothing or over-parameterization, and since training can be computationally demanding.
In order to address these shortcomings recent efforts have concentrated on finding simplified architectures that are both more interpretable and easier to optimize. Approaches to simplify GNNs have almost exclusively taken the point of view of node and graph classification. With our work, we aim to complement existing works by analysing simplified and untrained architectures from the perspective of link prediction Zhang and Chen (2018). Link prediction is an important task for graph learning algorithms with applications such as recommender systems, spam mail detection, drug repurposing, and many more Zhang and Chen (2018). Moreover, we show that link prediction can also provide a complementary perspective for the theoretical analysis of GNNs Zhou et al. (2022); Chamberlain et al. (2022).
@@
In this work, we explore the use of untrained message passing layers for link prediction in data on unlabelled graphs with high dimensional features. By formulating Untrained Message Passing (UTMP) layers, we follow an approach similar to that of Simplified Graph Convolutional Networks introduced by Wu et al.Wu et al. (2019). This approach simplifies GNN architectures by removing trainable parameters and nonlinearities resulting in an architecture that can be clearly separated into two components: an untrained message passing/feature propagation steps followed by a linear classifier. In addition to these we also consider fully untrained architectures based on simple inner products of features obtained after $l$ iterations of UTMP layers as a baseline and find that these features produced by UTMP layers are already highly informative leading to surprisingly high link prediction performances.
@@
We base our analysis on untrained versions of four widely used MP architectures, namely Graph Convolutional Networks (GCN)Kipf and Welling (2016a), SAGE Hamilton et al. (2017), GraphConv Morris et al. (2019) and GIN Xu et al. (2018). We test these untrained message passing layers on a variety of datasets that cover a wide range of sizes, node features, and topological characteristics ensuring a comprehensive evaluation of the models. Interestingly, for the large majority of datasets, we find that untrained message passing layers actually perform better than their fully trained counterparts while being highly interpretable and much easier to optimize.
@@
In our theoretical analysis we establish a direct connection between untrained message passing layers and path based measures. Path based methods and measures can capture the indirect connection strength between node pairs nodes in the absence of a direct link connecting the nodes. Consequently, path based measures and methods have been widely used in traditional link prediction methods Martínez et al. (2016); Kumar et al. (2020) and also play a key role in the state of the art methods Zhu et al. (2021); Zhang and Chen (2018). Our theoretical analysis is based on the assumption that initial node features are orthonormal which covers widely used initialization schemes such as as one-hot encodings and high dimensional random features, and also holds approximately for many empirical data sets with high dimensional features. Hence, our theoretical findings also provide new insights into the effectiveness of the widely used initialization schemes of one-hot encodings and high dimensional random features in graph representation learning. Our results show that untrained versions of message passing layers are highly amenable to theoretical analysis and hence could potentially serve as an general ansatz for the theoretical analysis of GNNs in settings beyond link prediction.
We have structured the remainder of the paper as follows: In Section 2, we summarize related works on untrained and simplified GNNs. Section 3 introduces the untrained message passing architectures that form the basis of our investigation. Here, we further provide a theoretical analysis of these layers from the perspective of node similarity measures that account both for node features and the graph topology. Section 4 details the experimental setup and results. Finally, in Section 5 we present our conclusions and highlight the potential for future work.
To ensure replicability of our results, we make our code available online  [^1] https://zenodo.org/records/11237762.
@@ 2 Related Work
Our work is motivated by recent works that investigate simplified and untrained GNNs from different perspectives. We follow an approach that is closely related to Wu et al. (2019) which simplifies GNNs by successively removing trainable parameters from layers and nonlinearities between consecutive layers. This results in an architecture that can be separated into two components. The first component consists of feature propagation in untrained message passing layers of which the output is then fed into a linear layer resulting in models that scale to larger data sets while being naturally interpretable. Wu et al. also provide a theoretical analysis of simplified models in the context of node classification reducing the model to a fixed low-pass filter followed by a linear classifier. The paper also empirically evaluates the simplified architectures on various downstream applications and shows that simplified architectures do not negatively impact accuracy while being computationally more efficient than their more complex counterparts.
@@
Other works have focused on finding untrained subnetworks. For instance Huang et al. (2022) explores the existence of untrained subnetworks in GNNs that can match the performance of fully trained dense networks at initialization, without any optimization of the weights. The paper leverages sparsity as the core tool to find such subnetworks and shows that they can substantially mitigate the over-smoothing problem, hence enabling deeper GNNs without bells and whistles. The paper also shows that the sparse untrained subnetworks have appealing performance in out-of-distribution detection and robustness to input perturbations. Similarly, in Böker et al. (2023) the authors demonstrate that GNNs with randomly initialized weights, without training, can achieve competitive performance compared to their trained counterparts focusing on the problem of graph classification. Other more recent works on untrained GNNs include Dong et al. (2024) where the authors propose a training free linear GNN model for semi supervised node classification in text attributed graphs and Sato (2024) that defines training free GNNs for transductive node classification based on using training labels as features.
@@
Link prediction is widely studied problem with a multitude of methods. Our method is related to a both GNN based methods where the prevalent approach is the use of Variational Graph Autoencoders (V-GAE) Kipf and Welling (2016b) and more traditional methods that are rely on path and random walk based measures for link prediction Kumar et al. (2020). Indeed, one of our main results is to establish a link between GNN based approaches and traditional path based measures. Some more recent methods such as SEAL Zhang and Chen (2018) and WalkPool Pan et al. (2021) rely on extracting local subgraphs around target links and frame link prediction as a graph classification problem. Although subgraph extraction based methods can out perform GNN based methods in link prediction tasks the subgraph extraction process can be very resource intensive for large networks negatively affecting the scalability of these methods. NBFNet Zhu et al. (2021) is another state-of-the-art method that is motivated by the Bellman-Ford algorithm. NBFNet is based on learning representations of paths between target nodes and aggregation functions for these representations. While NBFNet scales more favorably compared to subgraph extraction based methods it still needs to compute representations for large numbers of paths to predict links and hence has worse scaling behavior compared to purely GNN based link prediction methods Zhu et al. (2021).
@@
From a theoretical perspective, our results also relate to recent works on the effectiveness of random node initializations and one-hot encodings. For instance Sato et al. (2021) and Abboud et al. (2020) focus on the effect of random node initializations of the expressivity of GNNs in the context of graph classification while Cui et al. (2022) explore various encodings for the task including node and graph classification. Our analysis complements these works from the point of view of link classification and establishes a link between features derived from random and one hot initializations and, path based topological features.
@@ 3 Message Passing Architectures
Prior to introducing the message passing architectures investigated in our work, we first clarify the notation used throughout the paper. Let $G(V,E)$ be an undirected graph with vertex set $V=\{v_{1},v_{2}\ldots v_{N}\}$, edges $E\subseteq V\times V$ and no self-loops, i.e. $(v,v)\notin E\quad\forall v\in V$. We denote the adjacency matrix of the graph as $A$ and define $\tilde{\mathbf{A}}:=\mathbf{A}+\mathbf{I}$, where $\mathbf{I}$ is the identity matrix, i.e. $\tilde{\mathbf{A}}$ denotes the adjacency matrix of the graph $G$ that explicitly includes all self-loops. We use $\mathcal{N}(v)$ to denote the neighborhood of a node $v$, i.e. the set $\{w\in V:(v,w)\in E\}$, and use $\tilde{\mathcal{N}}(v):=\mathcal{N}(v)\cup\{v\}$ to denote the neighborhood of $v$ in the graph with self-loops. Similarly, we denote the degree of a node $v$ as $d(v)$ and $\tilde{d}(v)=d(v)+1$. The initial feature vector of node $v$ is denoted as $h_{v}^{(0)}$ and we use $h_{v}^{(l)}$ to denote the updated feature vector of node $v$ after $l$ rounds of message passing. Although we restrict our discussion to undirected and unweighted graphs the generalization of our definitions and results to weighted graphs is straightforward.
Prior to defining untrained versions, we first introduce the message passing rules of the four GNN architectures considered in our work, using the unified notation above.
Graph Convolutional Networks Kipf and Welling (2016a)
@@
$h_{v}^{(l)}=W^{(l),\top}\sum_{u\in\tilde{\mathcal{N}}(v)}\frac{1}{\sqrt{\tilde{d}_{u}\tilde{d}_{v}}}h_{u}^{(l-1)}$
where $W^{(l),\top}$ is the transposed weight matrix for this layer.
GraphSAGE Hamilton et al. (2017)
@@
$h_{v}^{(l)}=W^{(l)}_{1}h_{v}^{(l-1)}+W^{(l)}_{2}\cdot\text{AGG}_{u\in\mathcal{N}(v)}h_{u}^{(l-1)},$
where $W^{(l)}_{\{1,2\}}$ are learned weight matrices at layer $l$, $AGG$ is an aggregation function (such as mean, sum, max).
Throughout this paper use the following slightly modified version of the SAGE layer :
$h_{v}^{(l)}=W_{1}^{(l)}\cdot\frac{1}{\tilde{d}_{v}}\sum_{u\in\tilde{\mathcal{N}}(v)}h_{u}^{(l-1)},$
which we found to produce superior results for link prediction.
GIN Xu et al. (2018)
The Graph Isomorphism Network Convolution (GIN) is a simple architecture that is provably the most expressive among the class of GNNs and is as powerful as the Weisfeiler-Lehman graph isomorphism test. It is designed to capture the local structure of a graph by aggregating feature information from a node’s neighborhood. The mathematical formula for GIN is as follows:
@@
where $\Theta$ denotes a Multilayer Perceptron after each message passing layer, which in our implementation includes two Linear layers and a Rectified Linear Unit (ReLU) activation function following each Linear layer (code adapted from Böker et al. (2023); Morris and Ningyuan Huang (2023)).
GraphConv Morris et al. (2019)
@@ 3.1 Untrained MP Architectures
For the purpose of our theoretical and experimental evaluation, we now define the untrained counterparts of the four Message Passing Neural Network (MPNN) architectures introduced in the previous section. We specifically eliminate all learnable components and replace them with identity matrices. Our objective is to obtain the simplest form for the update function that retains the general message passing strategy, which includes the predefined update message passing functions and aggregation methods while removing all learnable parameters and non-linearities. We obtain the following functions that capture the aggregation and update step in the untrained versions of the message passing layers:

### v2

@@ Abstract
In this work, we explore the use of untrained message passing layers in graph neural networks for link prediction. The untrained message passing layers we consider are derived from widely used graph neural network architectures by removing trainable parameters and nonlinearities in their respective message passing layers. Experimentally, we find that untrained message passing layers can lead to competitive and even superior link prediction performance compared to fully trained message passing layers while being more efficient, especially in the presence of high-dimensional features. We also provide a theoretical analysis of untrained message passing layers in the context of link prediction and show that the inner product of features produced by untrained message passing layers relate to common neighbour and path-based topological measures which are widely used for link prediction. As such, untrained message passing layers offer a more efficient alternative to trained message passing layers in link prediction tasks, with clearer theoretical links to classical path-based heuristics.
@@ 1 Introduction
Graph neural networks (GNNs) are a powerful class of machine learning models that can learn from graph-structured data, such as social networks, molecular graphs, and knowledge graphs. GNNs have emerged as an important tool in the machine learning landscape, due to their ability to model complex relationships and dependencies within data and have found applications in a variety of fields where data exhibits a complex topology that can be captured as a graph. This is shown by a multitude of studies, including [1, 2, 3, 4, 5], which highlight the versatility and adaptability of GNNs for machine learning tasks across a range of fields.
@@
One of the key concepts underlying GNNs is Message Passing (MP) [6], which operates by propagating and aggregating information between nodes in the graph, using message and update functions possibly with learnable parameters. However, designing effective GNNs can be challenging, as they can suffer from issues such as over-smoothing or over-parameterization, and because training GNNs can be computationally demanding. In order to address these shortcomings recent efforts have concentrated on finding simplified architectures that are both more interpretable and easier to optimize.
With our work, we aim to complement existing works on simplified and untrained MP architectures, previously formulated in the context of node classification, from the perspective of link prediction [7]. Link prediction (LP) is an important task in graph ML with many applications such as recommender systems, spam mail detection, drug re-purposing, and many more [7]. Current state-of-the-art LP methods [8] in general rely on a combination of GNNs and structural features; hence formulating effective and efficient MP architectures has the potential to further improve current LP methods in these regards. Therefore, we focus our analysis on MP layers rather than trying to formulate a novel end-to-end LP method.
@@
In this work, we explore the use of untrained message passing layers for link prediction in graph datasets with high dimensional features. By formulating Untrained Message Passing (UTMP) layers, we follow an approach similar to that of Simplified Graph Convolutional Networks introduced by [9]. This approach simplifies GNN architectures by removing trainable parameters and nonlinearities resulting in an architecture that can be clearly separated into two components: an untrained message passing/feature propagation steps followed by a linear classifier. In addition to these we also consider fully untrained architectures based on simple inner products of features obtained after $l$ iterations of UTMP layers as a baseline and find that these features produced by UTMP layers are already highly informative leading to surprisingly high link prediction performances.
@@
We base our analysis on untrained versions of four widely used MP architectures, namely Graph Convolutional Networks (GCN)[10], SAGE [11], GraphConv [12] and GIN [13]. We test these untrained message passing layers on a variety of datasets that cover a wide range of sizes, node features, and topological characteristics ensuring a comprehensive evaluation of the models. We test UTMP layers on a variety of datasets that cover a wide range of sizes, node features, and topological characteristics ensuring a comprehensive evaluation of the UTMP layers. Mirroring the results reported by [9] for node classification we find that UTMP layers in many cases outperform their fully trained counterparts in LP tasks, while being easier to optimize and exhibiting clearer theoretical connections to classical path-based measures.
We also show that link prediction provides a complementary perspective for the theoretical analysis of UTMP layers [14, 15]. In our theoretical analysis we establish a direct connection between features produced by UTMP layers and various path based node similarity measures. Path based measures capture the indirect connection strength between node pairs nodes in the absence of a direct link connecting the nodes. Consequently, path based measures and methods have been widely used in traditional link prediction methods [16, 17] and also play a key role in many state of the art methods [18, 7].
@@
Our theoretical analysis relies on the assumption that initial node features are orthonormal which covers widely used initialization schemes such as as one-hot encodings and high dimensional random features, and also holds approximately for many empirical data sets with high dimensional features. Hence, our theoretical findings also provide new insights into the effectiveness of the widely used initialization schemes of one-hot encodings and high dimensional random features in graph representation learning. More generally our results show that untrained versions of message passing layers are highly amenable to theoretical analysis and hence could potentially serve as a general ansatz for the theoretical analysis of GNNs including settings beyond link prediction.
The main contributions of the paper are as follows:
  1. 1.
    We show that untrained versions of widely used MP layers often outperform their fully trained counterparts in LP tasks.
  2. 2.
    We establish a direct connection between MPNNs and path based node similarity measures both of which are widely used in LP methods.
  3. 3.
    Our theoretical analysis further provides insights in to the effectiveness of widely used node initialization schemes such as one-hot-encodings and random features in graph neural networks.
@@ 2 Related Work
Our work is motivated by recent works that investigate simplified and untrained GNNs from different perspectives. We formulate UTMP layers following the approach of [9] which simplifies GNNs by successively removing trainable parameters from layers and nonlinearities between consecutive layers. [9] also provide a theoretical analysis of simplified models in the context of node classification reducing the model to a fixed low-pass filter followed by a linear classifier. The paper also empirically evaluates the simplified architectures on various downstream applications and shows that simplified architectures do not negatively impact accuracy while being computationally more efficient than their fully trained counterparts.
@@
Other works have focused on finding untrained subnetworks. For instance [19] explores the existence of untrained subnetworks in GNNs that can match the performance of fully trained dense networks at initialization, without any optimization of the weights. The paper leverages sparsity as the core tool to find such subnetworks and shows that they can substantially mitigate the over-smoothing problem, hence enabling deeper GNNs. The paper also shows that the sparse untrained subnetworks have appealing performance in out-of-distribution detection and robustness to input perturbations. Similarly, in [20] the authors demonstrate that GNNs with randomly initialized weights, without training, can achieve competitive performance compared to their trained counterparts focusing on the problem of graph classification. In [21] the authors show that certain common neighbour measures can be approximated by MPNNs initialised with random weights and node features without training. Other more recent works on untrained GNNs include [22] where the authors propose a training free linear GNN model for semi supervised node classification in text attributed graphs and [23] that defines training free GNNs for transductive node classification based on using training labels as features.
@@
Link prediction is widely studied problem with a multitude of available methods. UTMP layers are related to a both GNN based methods such as Variational Graph Autoencoders (V-GAE) [24] and more traditional methods that are rely on path and random walk based measures for link prediction [17]. On the other hand state of the art methods such as SEAL [7], NBFnet [18], BUDDY [15], Neo-GNN [25] and NCNC [8] in general rely on combining GNNs and structural features. Some methods such as SEAL [7] and WalkPool [26] are based on extracting and performing MP on local subgraphs around target links effectively framing link prediction as a graph classification problem. Although subgraph extraction based methods can out perform purely GNN based methods in link prediction tasks the subgraph extraction process can be resource intensive for large networks negatively affecting the scalability of these methods. NBFnet [18] is another state-of-the-art method that is motivated by the Bellman-Ford algorithm. NBFnet is based on learning representations of paths between target nodes and aggregation functions for these representations. While NBFnet scales more favorably compared to subgraph extraction based methods it still needs to compute representations for large numbers of paths to predict links and hence has worse scaling behavior compared to purely GNN based link prediction methods [18]. Neo-GNN [25] and BUDDY [15] circumvent these difficulties by using pairwise similarity measures between higher order neighbourhoods of nodes, which are then used together with GNN based node features for LP. Notably BUDDY includes a feature propagation step that can be seen as a special case of UTMP (see UTSAGE in Section 3.1). In [8] the authors propose a GNN based approach that in addition to using node level features also aggregates features of common neighbours for link prediction and further propose Neural Common Neighbour Completion (NCNC) to counteract the negative effects of graph incompleteness on LP performance.
As detailed above GNNs are widely employed as sub-components in LP methods. Hence rather than proposing a new method for link prediction we consider the advantages of using untrained MP layers over their trained counterparts in existing LP methods. Moreover, our theoretical results establish a link between MP based approaches and structural features by showing that features resulting from UTMP implicitly encode neighbourhood information that underlies many widely used common neighbour and path based structural features used in LP.
@@
From a theoretical perspective, our results also relate to recent works on the effectiveness of random node initializations and one-hot encodings. For instance [27] and [28] focus on the effect of random node initializations of the expressivity of GNNs in the context of graph classification while [29] explores various encodings for the task including node and graph classification. Our analysis complements these works from the point of view of LP and establishes a link between features derived from random and one hot initializations and path based topological features. In addition, complementary work on initialization strategies such as the graph partition based GPA initializer [30] investigates how better initialization can improve downstream link prediction and node classification without changing the model architecture.
Finally, in our theoretical analysis we rely on the fact that collections of high dimensional vectors tend to be mutually orthogonal. This is a widely known fact that has wide-ranging applications in ML more broadly given the pervasive use of high dimensional vector representations in modern ML methods [31, 32].
@@ 3 Message Passing Architectures
Prior to introducing the message passing architectures investigated in our work, we first clarify the notation used throughout the paper. Let $G(V,E)$ be an undirected graph with vertex set $V=\{v_{1},v_{2}\ldots v_{N}\}$, edges $E\subseteq V\times V$ and no self-loops, i.e. $(v,v)\notin E\quad\forall v\in V$. We denote the adjacency matrix of the graph as $A$ and define $\tilde{\mathbf{A}}:=\mathbf{A}+\mathbf{I}$, where $\mathbf{I}$ is the identity matrix, i.e. $\tilde{\mathbf{A}}$ denotes the adjacency matrix of the graph $G$ that explicitly includes all possible self-loops. We use $\mathcal{N}(v)$ to denote the neighborhood of a node $v$, i.e. the set $\{w\in V:(v,w)\in E\}$, and use $\tilde{\mathcal{N}}(v):=\mathcal{N}(v)\cup\{v\}$ to denote the neighborhood of $v$ in the graph with self-loops. Similarly, we denote the degree of a node $v$ as $d(v)$ and $\tilde{d}(v)=d(v)+1$. The initial feature vector of node $v$ is denoted as $h_{v}^{(0)}$ and we use $h_{v}^{(l)}$ to denote the updated feature vector of node $v$ after $l$ rounds of message passing. Although we restrict our discussion to undirected and unweighted graphs the generalization of our definitions and results to weighted graphs is straightforward.
Prior to defining untrained versions, we first introduce the message passing rules of the four GNN architectures considered in our work, using a unified notation above.
On the inclusion of self-loops
Throughout our formulations we explicitly include self-loops, i.e., work with $\tilde{\mathcal{N}}(v)=\mathcal{N}(v)\cup\{v\}$ and $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$. This choice is motivated by both theoretical and practical considerations. First, self-loops allow nodes to retain their own features during aggregation, which prevents a complete dilution of node-specific information as the number of layers increases. Second, self-loops enable UTMP layers to capture paths of all lengths up to $2l$ (rather than exactly length $2l$), which makes the connection to path-based measures such as Katz index and rooted PageRank more direct. Third, this aligns our untrained formulations with standard trained architectures (e.g., GCN, SAGE, GraphConv), ensuring a fair comparison when removing trainable parameters.
Graph Convolutional Networks [10]
@@
$h_{v}^{(l)}=W^{(l)}\cdot\sum_{u\in\tilde{\mathcal{N}}(v)}\frac{1}{\sqrt{\tilde{d}_{u}\tilde{d}_{v}}}h_{u}^{(l-1)},$
where $W^{(l)}$ is the weight matrix for layer $l$.
GraphSAGE [11]
@@
$h_{v}^{(l)}=W^{(l)}_{1}\cdot h_{v}^{(l-1)}+W^{(l)}_{2}\cdot\text{AGG}_{u\in\mathcal{N}(v)}h_{u}^{(l-1)},$
where $W^{(l)}_{\{1,2\}}$ are learned weight matrices at layer $l$, AGGAGG is an aggregation function (such as mean, sum, max).
Throughout this paper use the following slightly modified version of the SAGE layer : $h_{v}^{(l)}=W_{1}^{(l)}\cdot\frac{1}{\tilde{d}_{v}}\sum_{u\in\tilde{\mathcal{N}}(v)}h_{u}^{(l-1)},$ which we found to produce superior results for link prediction.
GIN [13]
The Graph Isomorphism Network Convolution (GIN) is a simple architecture that is provably the most expressive among the class of GNNs and is as powerful as the Weisfeiler-Lehman graph isomorphism test. The MP step of GIN is as follows:
@@
where $\Theta$ denotes an MLP after each message passing layer, which in our implementation includes two Linear layers and a Rectified Linear Unit (ReLU) activation function following each Linear layer (code adapted from [20, 33]).
GraphConv [12]
@@ 3.1 Untrained MP Architectures
For the purpose of our theoretical and experimental evaluation, we now define the untrained counterparts of the four Message Passing Neural Network (MPNN) architectures introduced in the previous section. Following [9] we eliminate all learnable components and replace them with identity matrices. Here our objective is to obtain the simplest form for the update function that retains the general message passing strategy, which includes the predefined update message passing functions and aggregation methods while removing all learnable parameters and nonlinearities. We obtain the following functions that capture the aggregation and update step in the untrained versions of the message passing layers: